# Point2Primitive: CAD Reconstruction from Point Cloud by Direct Primitive Prediction

### v2

@@ 5.6 Ablation Study
We perform ablation studies to carefully analyze the DeNoising (DN), the improved transformer decoder, the sketch primitive definition, and the balancing weight in the loss function. All quantitative metrics are measured on the DeepCAD dataset.
No. $\beta$ SkhRep ImpDec $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$
@@
Table 5: Ablation study. left panel is the accuracy curves during training; right panel is the quantitative results.
DeNoising. It can be seen from the convergence curve shown in Figure 5 that by employing DeNoising in the training procedure, the ultimate sketch prediction accuracy is improved by $11.4\%$ and $3.1\%$ on $Acc^{SKH}_{t}$ and $Acc^{SKH}_{p}$, respectively. More validation loss curves during training are presented in Appendix 5. From the curves, we can see that by directly predicting the curve parameters and types, the accuracy is extremely high (type error below 0.5% and parameter error below 0.06)
Decoder and Sketch Hierarchy. We replace the improved decoder layer (ImpDec) with the vanilla transformer. Also, we utilize the sketch representation (SkhRep) method in [37], which converts parameter prediction into a classification problem. As shown in Table 5, the $Acc^{SKH}_{t}$ improves by $12.2\%$ by utilizing the ImpDec. Also, the center-prior sketch hierarchy contributes to an $8.4\%$ improvement in $Acc_{p}$. This shows the effectiveness of the proposed improved transformer decoder.
Balancing Weight. Table 5 shows that the balance weight $\beta=2$ performs slightly better than the other conditions, and we set it to our default value.
@@ 6 Conclusion
Supplementary Material
1 Evaluation Metrics
1.1 Primitive Type and Parameter Metrics
@@
$ACC_{type}=\frac{1}{N_{c}}\sum\limits_{i=1}^{N_{c}}\Psi[t_{i}=\hat{t}_{i}]$ (7)
@@
$ACC_{param}=\frac{1}{K}\sum\limits_{i=1}^{N_{c}}\sum\limits_{j=1}^{\lvert\hat{p}_{i}\rvert}\Phi[\lvert p_{i,j}-\hat{p}_{i,j}\rvert<\epsilon]\Psi[C_{i}=\hat{C}_{i}]$ (8)
@@ 1.2 Extrusion Segmentation IoU
$\mathrm{Seg\ IoU}$ $=\frac{1}{K}\sum_{k=1}^{K}RIoU(\mathbbm{1}(\mathbf{\hat{W}}_{:,k}),\mathbf{W}_{:,k})$ (9)
$RIoU(\mathbf{X},\mathbf{Y})$ $=\frac{\mathbf{X}^{\top}\cdot{\mathbf{Y}}}{{\left\|\mathbf{X}\right\|}_{1}+{\left\|\mathbf{Y}\right\|}_{1}-\mathbf{X}^{\top}\cdot{\mathbf{Y}}}$ (10)
@@
1.3 Base/Barrel Point Segmentation
@@
$\mathrm{BB\ Acc}=\frac{1}{N}\sum_{i=1}^{N}(\mathbbm{1}(\mathbf{\hat{B}}_{i,:})==\mathbf{B}_{i,:})$ (11)
@@
1.4 Extrusion Axis Angle Error
@@
$\mathrm{EA\ Angle\ Error}=\frac{1}{K}\sum_{k=1}^{K}\arccos(\mathbf{\hat{e}}^{\top}_{k,:}\mathbf{e}_{k,:})$ (12)
@@
1.5 Extrusion Fitting Error
@@
$\mathrm{Fit\ Ext}$ $=\frac{1}{K}\sum_{k=1}^{K}\mathcal{F}$ (13)
$\mathcal{F}$ $\triangleq\frac{1}{N_{\mathbf{P}_{k}}}\sum\left|SDF(\hat{\mathbf{s}}_{k},\hat{\mathbf{S}}_{k})-SDF(\mathbf{s}_{k},\mathbf{S}_{k}))\right|$ (14)
$\hat{\mathbf{S}}_{k}$ $=\prod\limits(\mathbf{P}^{barrel_{k}},\hat{e}_{k,:},\hat{c}_{k,:})$ (15)
where $\hat{c}_{k,:}$ is the predicted extrusion center. ($\prod$($\cdot$)∏ ( ⋅ ) projects per extrusion barrel points using the extrusion axis and center. The inner summation $\mathcal{F}(\mathbf{P},k)$ represents the goodness of the $k$-th extrusion fitting.
1.6 Primitive Number
@@
$\Delta\#P=\left|\hat{N}_{c}-N_{c}\right|$ (16)
@@
2 Visualized Sketch Prediction
@@
We provide more visualized results of the predicted sketch. Fig. 7 demonstrates some examples of the predicted sketch primitives. It can be seen that the predicted sketches are accurate both in curve type and parameter. However, some parameter errors can be seen in the Figure. Therefore, we develop a post-optimized method to eliminate the parameter errors, which is detailed in Section 5.2.
3 Validation Loss w/o DeNoising
Figure 8: Validation curves with and without DeNoising of the Point2Primitive during training: (a) the class error during training; (b) the L-1 loss of the primitive parameter during training; (c) the distance loss of the primitive parameter during training.
@@ 3 Validation Loss w/o DeNoising
4 The Implementation of the Extrusion Segmentation Network
@@
4.1 Extrusion and Barrel Point Segmentaion
@@
$\mathbf{\hat{W}}_{:,j}$ $=\mathbf{\hat{M}}_{:,2j}+\mathbf{\hat{M}}_{:,2j+1}$ (17)
$\mathbf{\hat{B}}_{:,0}$ $=\sum_{i}\mathbf{\hat{M}}_{:,2j}$ (18)
$\mathbf{\hat{B}}_{:,1}$ $=\sum_{i}\mathbf{\hat{M}}_{:,2j+1}$ (19)
4.2 Sketch Point Encoder
As shown in Figure 9, the sketch point encoder is developed based on the PointCNN[16]. Instead of using random or fps points, fixed points are set as the representation points. Also, to provide position encodings for the transformer, a learned position encoding is utilized.
5 Post Optimization
5.1 The Batch SDF Computing
@@
5.2 Curve Fine-tuning
@@
$Curves_{norm}\leftarrow sort(Curves_{norm})$triangleright sorted by starting point
@@
$T\leftarrow\emptyset$triangleright as modified Curves
@@ 5.3 Loop Opmitzation
$point\_cloud,Curves$
@@
$eval\_points\leftarrow in\_pc(point\_cloud)$triangleright as pc to be formulated
$cmd,params\leftarrow Initialize(sketch)$triangleright cmd as mask of curve type, params as shape of sketch
for each epoch in $epoches$ do
       $line\_loss\leftarrow SDF(eval\_points,params[\newline line\_mask(cmd)])$triangleright calculate line, arc and circle loss
       $arc\_loss\leftarrow SDF(eval\_points,params[\newline arc\_mask(cmd)])$
       $circle\_loss\leftarrow SDF(eval\_points,params[\newline circle\_mask(cmd)])$
       $loss\leftarrow concat(line\_loss,arc\_loss,circle\_loss,-1)$
       $loss.min()$triangleright minimize the last element of loss tuple
@@
6 Extrusion Parameters Calculation
We calculate the parameters of the extrusion operation by the obtained barrel points. The parameters of the extrusion operation $E_{i}=(\hat{e_{i}},\hat{t_{i}})$ consists of the extrusion axis $\hat{e_{i}}$, extrusion position $\hat{o_{i}}$, and extrusion extent $\hat{t_{i}}$. Firstly, the optimal extrusion axis of an input extrusion cylinder PC is given by $\hat{e_{i}}=\mbox{argmin}_{e_{i},\|e_{i}\|=1}(e_{i}^{T}H_{\phi}e_{i})$, where:
$H_{\phi}=\textbf{N}^{\top}\Phi_{barr}^{\top}\Phi_{barr}\textbf{N}-\textbf{N}^{\top}\Phi_{base}^{\top}\Phi_{base}\textbf{N}$ (20)
@@ 6 Extrusion Parameters Calculation
$\begin{split}&\hat{t}_{min}=\min\limits_{P_{i}\in{P_{barr}}}(e_{i}\cdot(P_{i}-\hat{o}))\\ &\hat{t}_{max}=\max\limits_{P_{i}\in{P_{barr}}}(e_{i}\cdot(P_{i}-\hat{o}))\\ \end{split}$ (21)
7 More Details of the Improved Transformer Decoder
7.1 Command Parameter Noise
@@ 7.2 Attention Mask
$\begin{split}&q=\{g_{0},g_{1},...,g_{K-1}\}\\ &g_{j}=\{q_{0}^{k},q_{1}^{k},...,q_{M-1}^{k}\}\end{split}$ (22)
where $g_{k}$ is the $-k\mbox{-}$th denosing group. $q_{m}^{k}$ is the $-m\mbox{-}$th query in the denoising group. Each denoising group contains $M$ queries where $M$ is the number of commands in one input PC.
@@
$a_{ij}=\left\{\begin{aligned} 1&,\text{ if }j<K\times{M}\text{ and }\lfloor\frac{i}{M}\rfloor\neq\lfloor\frac{j}{M}\rfloor;\\ 1&,\text{ if }j<K\times{M}\text{ and }i\geq K\times{M};\\ 0&,\text{ else }\\ \end{aligned}\right.$ (23)
where $a_{ij}=1$ means the $-i\mbox{-}$th query cannot see the $-j\mbox{-}$th query and $a_{ij}=0$ other wise.
The decoder embeddings are taken as curve embeddings in our model. Therefore, an indicator is appended to the curve embeddings to distinguish the denoising part from the matching part, as shown in Figure 3, where $1$ means the query belongs to the denoising part and 00 means the query belongs to the matching part.
8 Applications

### v3

@@ 4.6 Ablation Study
We perform ablation studies to carefully analyze the query denoising (DN), the improved transformer decoder, the sketch primitive definition, and the balancing weight in the loss function. All quantitative metrics are measured on the DeepCAD dataset.
(a) Accuracy curve.
No. $\beta$ SkhRep ImpDec $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ CDCD
@@
(b) Quantitative results of ablation study.
Figure 6: Ablation study.
Query denoising. It can be seen from the convergence curve shown in Figure 6(a) that by employing denoising in the training procedure, the ultimate sketch prediction accuracy is improved by $11.4\%$ and $3.1\%$ on $Acc^{SKH}_{t}$ and $Acc^{SKH}_{p}$, respectively. Please refer to Appendix. D for more visualizations. From the curves, one can see that by directly predicting the curve parameters and types, the accuracy is extremely high (type error below 0.5% and parameter error below 0.06).
Decoder and Sketch Hierarchy. We replace the improved decoder layer (ImpDec) with the vanilla transformer. Also, we utilize the sketch representation (SkhRep) method in Wu_Xiao_Zheng_2021, which converts parameter prediction into a classification problem. As shown in Figure 6(b), the $Acc^{SKH}_{t}$ improves by $12.2\%$ by utilizing the ImpDec. Also, the center-prior sketch hierarchy contributes to an $8.4\%$ improvement in $Acc_{p}$. This shows the effectiveness of the proposed improved transformer decoder.
Balancing Weight. Figure 6(a) shows that the balance weight $\beta=2$ performs slightly better than the other conditions, and we set it to our default value.
@@ 5 Conclusion
Appendix A The Use of Large Language Models (LLMs)
In this work, Large Language Models (LLMs) were used solely to polish the language for clarity and readability. No LLMs were employed for idea generation, experimental design, data analysis, or any other part of the research process.
Appendix B Evaluation Metrics
B.1 Primitive Type and Parameter Metrics
@@
$ACC_{\text{type}}=\frac{1}{N_{c}}\sum\limits_{i=1}^{N_{c}}\Psi[t_{i}=\hat{t}_{i}],$ (9)
@@
$ACC_{\text{param}}=\frac{1}{K}\sum\limits_{i=1}^{N_{c}}\sum\limits_{j=1}^{\lvert\hat{p}_{i}\rvert}\Phi[\lvert p_{i,j}-\hat{p}_{i,j}\rvert<\epsilon]\Psi[C_{i}=\hat{C}_{i}],$ (10)
@@ B.2 Extrusion Segmentation IoU
$\mathrm{Seg\ IoU}$ $=\frac{1}{K}\sum_{k=1}^{K}RIoU(\mathbbm{1}(\mathbf{\hat{W}}_{:,k}),\mathbf{W}_{:,k}),$ (11)
$RIoU(\mathbf{X},\mathbf{Y})$ $=\frac{\mathbf{X}^{\top}\cdot{\mathbf{Y}}}{{\left\|\mathbf{X}\right\|}_{1}+{\left\|\mathbf{Y}\right\|}_{1}-\mathbf{X}^{\top}\cdot{\mathbf{Y}}},$ (12)
@@
B.3 Base/Barrel Point Segmentation
@@
$\mathrm{BB\ Acc}=\frac{1}{N}\sum_{i=1}^{N}(\mathbbm{1}(\mathbf{\hat{B}}_{i,:})==\mathbf{B}_{i,:}),$ (13)
@@
B.4 Extrusion Axis Angle Error
@@
$\mathrm{EA\ Angle\ Error}=\frac{1}{K}\sum_{k=1}^{K}\arccos(\mathbf{\hat{e}}^{\top}_{k,:}\mathbf{e}_{k,:}),$ (14)
@@
B.5 Extrusion Fitting Error
@@
$\mathrm{Fit\ Ext}$ $=\frac{1}{K}\sum_{k=1}^{K}\mathcal{F},\quad\text{where}$ (15)
$\mathcal{F}\triangleq\frac{1}{N_{\mathbf{P}_{k}}}\sum$ $\left|\texttt{SDF}(\hat{\mathbf{s}}_{k},\hat{\mathbf{S}}_{k})-\texttt{SDF}(\mathbf{s}_{k},\mathbf{S}_{k}))\right|,\quad\text{and}\quad\hat{\mathbf{S}}_{k}=\prod\limits(\mathbf{P}^{barrel_{k}},\hat{e}_{k,:},\hat{c}_{k,:}),$ (16)
where $\hat{c}_{k,:}$ is the predicted extrusion center. $\prod(\cdot)$ projects per extrusion barrel points using the extrusion axis and center. The inner summation $\mathcal{F}(\mathbf{P},k)$ represents the goodness of the $k$-th extrusion fitting.
B.6 Primitive Number
@@
$\Delta\#P=\left|\hat{N}_{c}-N_{c}\right|,$ (17)
@@
Appendix C Visualized Sketch Prediction
@@
We provide more visualized results of the predicted sketch. Fig. 7 demonstrates some examples of the predicted sketch primitives. It can be seen that the predicted sketches are accurate both in curve type and parameter. However, some parameter errors can be seen in the Figure. Therefore, we develop a post-optimized method to eliminate the parameter errors, which is detailed in Section F.2.
Appendix D Validation Loss w/o query denoising
Figure 8: Validation curves with and without query denoising of the Point2Primitive during training: (a) the class error during training; (b) the L-1 loss of the primitive parameter during training; (c) the distance loss of the primitive parameter during training.
@@ Appendix D Validation Loss w/o query denoising
Appendix E The Implementation of the Extrusion Segmentation Network
@@
E.1 Extrusion and Barrel Point Segmentaion
@@
$\hat{\mathbf{W}}_{:,j}$ $=\hat{\mathbf{M}}_{:,2j}+\hat{\mathbf{M}}_{:,2j+1},$ (18)
$\hat{\mathbf{B}}_{:,0}$ $=\sum_{j}\hat{\mathbf{M}}_{:,2j},\quad\text{and}\quad\hat{\mathbf{B}}_{:,1}=\sum_{j}\hat{\mathbf{M}}_{:,2j+1}.$ (19)
E.2 Sketch Point Encoder
As shown in Figure 9, the sketch point encoder is developed based on the PointCNN(li2018pointcnnconvolutionmathcalxtransformedpoints). Instead of using random or fps points, fixed points are set as the representation points. Also, to provide position encodings for the transformer, a learned position encoding is utilized.
Appendix F Post Optimization
F.1 The Batch SDF Computing
@@
F.2 Curve Fine-tuning
@@
$Curves_{norm}\leftarrow sort(Curves_{norm})$ $\triangleright$ sorted by starting point
@@
$T\leftarrow\emptyset$ $\triangleright$ as modified Curves
@@ F.3 Loop Opmitzation
point_cloud, Curves
@@
eval_points $\leftarrow$ in_pc(point_cloud) $\triangleright$ as pc to be formulated
cmd, params $\leftarrow$ Initialize(sketch) $\triangleright$ cmd as mask of curve type, params as shape of sketch
for each epoch in epochesepoches do
  line_loss $\leftarrow$ SDF(eval_points, params[line_mask(cmd)]) $\triangleright$ calculate line, arc and circle loss
  arc_loss $\leftarrow$ SDF(eval_points, params[arc_mask(cmd)])
  circle_loss $\leftarrow$ SDF(eval_points, params[circle_mask(cmd)])
  loss $\leftarrow$ concat(line_loss, arc_loss, circle_loss, -1)
  loss.min() $\triangleright$ minimize the last element of loss tuple
@@
Appendix G Extrusion Parameters Calculation
[…]
$H_{\phi}=\textbf{N}^{\top}\Phi_{barr}^{\top}\Phi_{barr}\textbf{N}-\textbf{N}^{\top}\Phi_{base}^{\top}\Phi_{base}\textbf{N},$ (20)
@@ Appendix G Extrusion Parameters Calculation
[…]
Appendix H More Details of the Improved Transformer Decoder
H.1 Command Parameter Noise
@@ H.2 Attention Mask
$\begin{split}q=\{g_{0},g_{1},...,g_{K-1}\},\quad\text{and}\quad g_{j}=\{q_{0}^{k},q_{1}^{k},...,q_{M-1}^{k}\}\end{split}$ (22)
where $g_{k}$ is the $k\mbox{-}$th denosing group. $q_{m}^{k}$ is the $m\mbox{-}$th query in the denoising group. Each denoising group contains $M$ queries where $M$ is the number of commands in one input PC.
@@
[…]
where $a_{ij}=1$ means the $i\mbox{-}$th query cannot see the $j\mbox{-}$th query and $a_{ij}=0$ other wise.
The decoder embeddings are taken as curve embeddings in our model. Therefore, an indicator is appended to the curve embeddings to distinguish the denoising part from the matching part, as shown in Figure 2, where $1$ means the query belongs to the denoising part and 0 means the query belongs to the matching part.
Appendix I More visualizations
We provide more visualization results as shown in Figure 11 and Figure 12.
Figure 11: More visual comparison between reconstruction results on the DeepCAD and Fusion360 gallery dataset.
Figure 12: More visual comparison between reconstruction results on the augmented DeepCAD dataset.
Appendix J Applications